OpenStudy (anonymous):

PRE-CAL: if log x=a, log y=b and log c=z, write log sqrt10xyz in terms of a,b, and c

OpenStudy (anonymous):

So, using the rules for logs, we can rewrite your logarithm. The specific rule I'm talking about is: \[Log(abc) = Log(a)+Log(c)+Log(d)\] First, use that to expand the log. You should end up with 4 different logs.

OpenStudy (anonymous):

And I have a clarification question. Are you given: \[Log(\sqrt{10xyz})\] or \[Log(\sqrt{10}xyz)\]

OpenStudy (anonymous):

yes, thats wat was given the first one

OpenStudy (ranga):

\[\Large \log(\sqrt{10xyz}) = \log(10xyz)^{\frac 12} = \frac 12 \log(10xyz) = \]

OpenStudy (ranga):

\[\Large \frac 12\{\log(10) + \log(x) + \log(y) + \log(z)\}\]

OpenStudy (ranga):

Put log(x) = a; log(y) = b; log(z) = c above.

OpenStudy (anonymous):

i'm a bit confuse with the last result u have set